OpenStudy (anonymous):

what are the solutions of the equation s^3+125=0

OpenStudy (anonymous):

evidently \(-5\) is one solution you need the two complex ones as well?

OpenStudy (ahsome):

Equation: \(s^3+125=0\) Move the 125 to the Right Hand Side \[s^3+125=0\]\[s^3=-125\]Now use the cube root:\[s=\sqrt[3]{125}\]\[s=5\]